# Gaussian Process-Based Prediction and Control of Hammerstein-Wiener Systems

### v1

@@ III-B Kernel Design and Posterior Prediction
$K(\boldsymbol{\eta}^{d},\boldsymbol{\eta}^{d})$ $=\left(\mathbb{I}\otimes\left[\Gamma_{1}\ \Gamma_{2}\ -\mathbb{I}\right]\right)K_{uy}(\boldsymbol{\eta}^{d})\left(\mathbb{I}\otimes\begin{bmatrix}\Gamma_{1}^{\top}\\ \Gamma_{2}^{\top}\\ -\mathbb{I}\end{bmatrix}\right),$ (22)
where $H_{u}:=\mathcal{H}_{L}(\mathbf{u}^{d})$, $H_{y}:=\mathcal{H}_{L}(\mathbf{y}^{d})$,
@@
$\mathbf{k}_{uy}$ $:=\begin{bmatrix}\mathbf{k}_{u}(\mathrm{vec}\left(H_{u}\right),\mathbf{u})&\mathbf{0}\\ \mathbf{0}&\mathbf{k}_{y}(\mathrm{vec}\left(H_{y}\right),\mathbf{y})\end{bmatrix},$
@@
$K_{uy}$ $:=\begin{bmatrix}\mathbf{k}_{u}(\mathrm{vec}\left(H_{u}\right),\mathrm{vec}\left(H_{u}\right))&\mathbf{0}\\ \mathbf{0}&\mathbf{k}_{y}(\mathrm{vec}\left(H_{y}\right),\mathrm{vec}\left(H_{y}\right))\end{bmatrix}.$
@@
$\left[\Gamma_{1}\ \Gamma_{2}\right]\mathrm{col}\left(\Psi(\mathbf{u}),\Phi(\mathbf{y}_{\text{p}})\right)-\Phi(\mathbf{y}_{\text{f}})\sim\mathcal{N}\left(m_{\text{p}}(\eta),k_{\text{p}}(\eta)\right),$
@@
whereas the true noise-free future output trajectory $\mathbf{y}_{f,0}$ satisfies $\left[\Gamma_{1}\ \Gamma_{2}\right]\mathrm{col}\left(\Psi(\mathbf{u}),\Phi(\mathbf{y}_{\text{p}})\right)-\Phi(\mathbf{y}_{\text{f,0}})=\left[\Gamma_{2}\ \mathbf{0}\right]\mathbf{e}$ according to (2). Therefore, the Gaussian process model characterizes probable trajectories of (1) in the sense that for any given input trajectory $\mathbf{u}$ and past output trajectory $\mathbf{y}_{\text{p}}$, $\mathbf{y}_{\text{f}}$ is the future output trajectory with projected prediction error $\Phi(\mathbf{y}_{f,0})-\Phi(\mathbf{y}_{\text{f}})\sim\mathcal{N}\left(m_{\text{p}}(\eta),\hat{k}_{\text{p}}(\eta)\right)$, where $\hat{k}_{\text{p}}(\eta)=k_{\text{p}}(\eta)+\mathrm{cov}\left(\left[\Gamma_{2}\ \mathbf{0}\right]\mathbf{e}\right)=k_{\text{p}}(\eta)+\sigma^{2}\Gamma_{2}\Gamma_{2}^{\top}$.
Different optimality criteria can be used to obtain the optimal prediction.
1) Maximum likelihood method: it finds the $\mathbf{y}_{\text{f}}$ that maximizes the probability of observing $f(\eta)=\mathbf{0}$ given $\boldsymbol{\chi}^{d}=\mathbf{0}$, i.e.,
$\mathbf{y}_{\text{f}}$ $=\text{arg}\underset{\mathbf{y}_{\text{f}}}{\text{max}}\ p\left(f(\eta)=\mathbf{0}|\boldsymbol{\chi}^{d}=\mathbf{0}\right)$ (23)
$=\text{arg}\underset{\mathbf{y}_{\text{f}}}{\text{min}}\ \mathrm{logdet}\left(k_{\text{p}}(\eta)\right)+m_{\text{p}}^{\top}(\eta)k_{\text{p}}^{-1}(\eta)m_{\text{p}}(\eta).$ (24)
2) Minimum mean-squared error predictor:
$\mathbf{y}_{\text{f}}=\text{arg}\underset{\mathbf{y}_{\text{f}}}{\text{min}}\ \text{tr}\left(\hat{k}_{\text{p}}(\eta)\right)+m_{\text{p}}^{\top}(\eta)m_{\text{p}}(\eta).$ (25)
3) Minimum-variance unbiased predictor
$\mathbf{y}_{\text{f}}=\text{arg}\underset{\mathbf{y}_{\text{f}}}{\text{min}}\ g\left(\hat{k}_{\text{p}}(\eta)\right)\text{ s.t. }m_{\text{p}}(\eta)=\mathbf{0},$ (26)
@@
where $g(\cdot)$ can be any optimality criteria, such as trace, determinant, or the largest eigenvalue, i.e., the A-, D-, and E- optimality criteria.
Remark 6
@@
For Hammerstein systems, $k_{\text{p}}(\eta)$ does not depend on $\mathbf{y}_{\text{f}}$ since $\mathbf{k}_{y}(\cdot,\cdot)=\mathbf{0}$ as mentioned in Remark 5. Therefore, all optimality criteria admit their minimum when $m_{\text{p}}(\eta)=\mathbf{0}$, i.e., $m(\eta)-\mathbf{k}^{\top}(\boldsymbol{\eta}^{d},\eta)\left(K+\bar{\Sigma}\right)^{-1}\mathbf{m}(\boldsymbol{\eta}^{d})=\mathbf{0}$. This leads to an explicit predictor by using (18):
$\mathbf{y}_{\text{f}}=\left[\Gamma_{1}\ \Gamma_{2}\right]\mathrm{col}\left(\mathbf{m}_{u}(\mathbf{u}),\mathbf{y}_{\text{p}}\right)-\mathbf{k}^{\top}(\boldsymbol{\eta}^{d},\eta)\left(K+\bar{\Sigma}\right)^{-1}\mathbf{m}(\boldsymbol{\eta}^{d}).$
III-C Hyperparameter Tuning

### v2

@@ 3.2 Structured Kernel Design
$K(\boldsymbol{\eta}^{d},\boldsymbol{\eta}^{d})$ $=\left(\mathbb{I}\otimes\Gamma_{1}\right)\mathbf{k}_{u}(\mathrm{vec}\left(H_{u}\right),\mathrm{vec}\left(H_{u}\right))\left(\mathbb{I}\otimes\Gamma_{1}^{\top}\right)$
$\ +\left(\mathbb{I}\otimes\bar{\Gamma}_{2}\right)\mathbf{k}_{y}(\mathrm{vec}\left(H_{y}\right),\mathrm{vec}\left(H_{y}\right))\left(\mathbb{I}\otimes\bar{\Gamma}_{2}^{\top}\right),$ (24)
where $H_{u}:=\mathcal{H}_{L}(\mathbf{u}^{d})$, $H_{y}:=\mathcal{H}_{L}(\mathbf{y}^{d})$. The detailed calculation is presented in Appendix A.
By applying the structured kernel design (22)-(24), the GP model (11)-(13) of the implicit function $f(\eta)$ can be obtained with prior mean and covariance functions (20), (21), noise model (19), and training data (18).
For later derivations, we also define the joint distribution of $\Phi(\mathbf{y})$ and $\boldsymbol{\chi}^{d}$:
[…]
where, similar to the derivation in Appendix A,
$\kappa_{y}(\boldsymbol{\eta}^{d},\mathbf{y}):=\left(\mathbb{I}\otimes\bar{\Gamma}_{2}\right)\mathbf{k}_{y}(\mathrm{vec}\left(H_{y}\right),\mathbf{y}).$ (26)
Similarly, we also define the joint distribution of $\Psi(\mathbf{u})$ and $\boldsymbol{\chi}^{d}$ with
$\kappa_{u}(\boldsymbol{\eta}^{d},\mathbf{u}):=\left(\mathbb{I}\otimes\Gamma_{1}\right)\mathbf{k}_{u}(\mathrm{vec}\left(H_{u}\right),\mathbf{u}).$ (27)
3.3 Incorporating Monotonicity by Expectation Propagation
The GP model (11)-(13), however, does not incorporate the monotonicity of the output nonlinearity $\phi(\cdot)$ as required in Assumption 1.3. Due to the unbounded nature of Gaussian processes, it is not possible to guarantee monotonicity for all posterior samples with certainty. Instead, since the derivative of a Gaussian process is still Gaussian, we follow the idea described in [36] and add virtual derivative points to the training data to promote functions that are monotonically increasing.
Remark 5.
We note that if the posterior mean estimate of $\phi(\cdot)$ (cf. Section 3.6) is not monotonic, the predictor remains well-defined, but it does not correspond to a well-defined Wiener system.
For simplicity of exposition, we present the case where $n_{y}=1$. The extension to the multi-output case is discussed in Remark 6. We would like to include information that $\phi^{\prime}(y_{\mathrm{m},i})>0$ for a grid of virtual derivative points $\mathbf{y}_{\mathrm{m}}:=\left(y_{\mathrm{m},i}\right)_{i=1}^{n_{m}}$ in the GP regression scheme. See [36, Section 3.1] for alternative strategies to place virtual derivative points. Let $\mathcal{M}_{i}$ be the event that $\phi^{\prime}(y_{\mathrm{m},i})>0$. Similar to the probit regression in classification [35, Chapter 3], we use the following probit likelihood to approximate the monotonicity constraint $\phi^{\prime}(y_{\mathrm{m},i})>0$ with a smooth function:
$\mathrm{Pr}\left(\mathcal{M}_{i}|\phi^{\prime}(y_{\mathrm{m},i})\right)\approx F_{\mathcal{N}}\left(\phi^{\prime}(y_{\mathrm{m},i})/\nu_{0}\right),$ (28)
where $F_{\mathcal{N}}(\cdot)$ is the cumulative distribution function of the unit Gaussian distribution and $\nu_{0}>0$ is a small constant such that $F_{\mathcal{N}}\left(x/\nu_{0}\right)$ approximates the step function as $\nu_{0}\rightarrow 0$. Following this approximation, we replace $\mathrm{Pr}\left(\mathcal{M}_{i}|\phi^{\prime}(y_{\mathrm{m},i})\right)$ by $F_{\mathcal{N}}\left(\phi^{\prime}(y_{\mathrm{m},i})/\nu_{0}\right)$ in what follows.
Then, from Bayes’ rule, we have
$p(\Phi^{\prime}(\mathbf{y}_{\mathrm{m}}),\mathbf{f}(\boldsymbol{\eta}^{d})|\mathcal{M},\boldsymbol{\chi}^{d})\propto p(\Phi^{\prime}(\mathbf{y}_{\mathrm{m}}),\mathbf{f}(\boldsymbol{\eta}^{d}))\\ p(\boldsymbol{\chi}^{d}|\mathbf{f}(\boldsymbol{\eta}^{d}))\prod_{i=1}^{n_{m}}F_{\mathcal{N}}\left(\phi^{\prime}(y_{\mathrm{m},i})/\nu_{0}\right),$ (29)
where $\Phi^{\prime}(\mathbf{y}_{\mathrm{m}}):=\mathrm{col}\left(\phi^{\prime}(y_{\mathrm{m},1}),\dots,\phi^{\prime}(y_{\mathrm{m},n_{m}})\right)$, $\mathcal{M}:=\cap_{i=1}^{n_{m}}\mathcal{M}_{i}$, and $\boldsymbol{\chi}^{d}|\mathbf{f}(\boldsymbol{\eta}^{d})\sim\mathcal{N}\left(\mathbf{f}(\boldsymbol{\eta}^{d}),\bar{\Sigma}\right)$. The joint distribution of $\mathbf{f}(\boldsymbol{\eta}^{d})$ and $\Phi^{\prime}(\mathbf{y}_{\mathrm{m}})$ is calculated as
@@ 3.3 Incorporating Monotonicity by Expectation Propagation
$\begin{bmatrix}\Phi^{\prime}(\mathbf{y}_{\mathrm{m}})\\ \mathbf{f}(\boldsymbol{\eta}^{d})\end{bmatrix}\sim\mathcal{N}\left(\underbrace{\begin{bmatrix}\mathbf{m}^{\prime}_{y}(\mathbf{y}_{\mathrm{m}})\\ \mathbf{m}(\boldsymbol{\eta}^{d})\end{bmatrix}}_{\mathbf{m}_{\mathrm{j}}},\underbrace{\begin{bmatrix}\mathbf{k}^{\prime\prime}_{y}(\mathbf{y}_{\mathrm{m}},\mathbf{y}_{\mathrm{m}})&{\kappa^{\prime}_{y}}^{\top}(\boldsymbol{\eta}^{d},\mathbf{y}_{\mathrm{m}})\\ \kappa^{\prime}_{y}(\boldsymbol{\eta}^{d},\mathbf{y}_{\mathrm{m}})&K(\boldsymbol{\eta}^{d},\boldsymbol{\eta}^{d})\end{bmatrix}}_{K_{\mathrm{j}}}\right),$
where $\mathbf{m}^{\prime}_{y}(\mathbf{y}_{\mathrm{m}})$, $\mathbf{k}^{\prime\prime}_{y}(\mathbf{y}_{\mathrm{m}},\mathbf{y}_{\mathrm{m}})$, $\kappa^{\prime}_{y}(\boldsymbol{\eta}^{d},\mathbf{y}_{\mathrm{m}})$ apply $m^{\prime}_{y}(y_{\mathrm{m},i})$, $\frac{\partial k_{y}(y_{\mathrm{m},i},y_{\mathrm{m},j})}{\partial y_{\mathrm{m},i}\partial y_{\mathrm{m},j}}$, $\frac{\partial\kappa_{y}(\boldsymbol{\eta}^{d},y_{\mathrm{m},i})}{\partial y_{\mathrm{m},i}}$ elementwise, respectively.
Unfortunately, the posterior distribution (29) is no longer Gaussian due to the terms $F_{\mathcal{N}}\left(\phi^{\prime}(y_{\mathrm{m},i})/\nu_{0}\right)$. For tractability, we apply the expectation propagation (EP) algorithm [35, Section 3.6] by approximating $F_{\mathcal{N}}\left(\phi^{\prime}(y_{\mathrm{m},i})/\nu_{0}\right)$ with an un-normalized Gaussian likelihood.
A numerically stable and efficient EP algorithm is presented in Algorithm 1. This algorithm adapts [35, Algorithm 3.5] for classification problems to the virtual derivative point method, using the posterior marginal moment equations from [36, Section 3]. The details of the EP algorithm are discussed in Appendix B.
Algorithm 1 Expectation propagation for the virtual derivative point method
1:Input: $\mathbf{m}_{\mathrm{j}}$, $K_{\mathrm{j}}$, $\bar{\Sigma}$
@@
2:Initialize: $\mathbf{S}:=\left(K_{\mathrm{j}}^{-1}+\mathrm{blkdiag}\left(\mathbf{0},\bar{\Sigma}^{-1}\right)\right)^{-1}$, $\boldsymbol{\mu}:=\mathbf{S}K_{\mathrm{j}}^{-1}\mathbf{m}_{\mathrm{j}}$, $\boldsymbol{\nu}:=\mathbf{0}$, $\boldsymbol{\tau}:=\mathbf{0}$
3:repeat
4:  for $i:=1$ to $n_{m}$ do
5:   $\tau_{-i}:=(\mathbf{S})_{i,i}^{-1}-\tau_{i}$, $\nu_{-i}:=(\mathbf{S})_{i,i}^{-1}\mu_{i}-\nu_{i}$
6:   $z_{i}:=\dfrac{\nu_{-i}}{\nu_{0}\tau_{-i}\sqrt{1+1/(\nu_{0}^{2}\tau_{-i})}}$
7:   $\hat{\mu}_{i}:=\dfrac{\nu_{-i}}{\tau_{-i}}+\dfrac{p_{\mathcal{N}}(z_{i})}{\nu_{0}\tau_{-i}F_{\mathcal{N}}(z_{i})\sqrt{1+1/(\nu_{0}^{2}\tau_{-i})}}$
8:
$\hat{\sigma}_{i}^{2}:=\dfrac{1}{\tau_{-i}}-\dfrac{p_{\mathcal{N}}(z_{i})}{\tau_{-i}^{2}F_{\mathcal{N}}(z_{i})(\nu_{0}^{2}+1/\tau_{-i})}\left(z_{i}+\dfrac{p_{\mathcal{N}}(z_{i})}{F_{\mathcal{N}}(z_{i})}\right)$
9:   $\Delta\tau:=\hat{\sigma}_{i}^{-2}-\tau_{-i}-\tau_{i}$
10:   $\tau_{i}:=\tau_{i}+\Delta\tau$, $\nu_{i}:=\hat{\sigma}_{i}^{-2}\hat{\mu}_{i}-\nu_{-i}$
11:   $\mathbf{S}:=\mathbf{S}-\left((\Delta\tau)^{-1}+(\mathbf{S})_{i,i}\right)^{-1}\mathbf{s}_{i}\mathbf{s}_{i}^{\top}$, where $\mathbf{s}_{i}$ is the $i$-th column of $\mathbf{S}$.
12:   $\boldsymbol{\mu}:=\mathbf{S}\mathrm{col}\left(\boldsymbol{\nu},\mathbf{0}\right)+\mathbf{S}K_{\mathrm{j}}^{-1}\mathbf{m}_{\mathrm{j}}$
13:  end for
14:until convergence
15:Output: $\tilde{\boldsymbol{\mu}}_{\mathrm{j}}:=\mathrm{col}\left(\nu_{1}/\tau_{1},\dots,\nu_{n_{m}}/\tau_{n_{m}},\mathbf{0}\right)$, $\tilde{\Sigma}_{\mathrm{j}}:=\mathrm{blkdiag}\left(1/\tau_{1},\dots,1/\tau_{n_{m}},\bar{\Sigma}\right)$.
Remark 6.
When $n_{y}>1$, the monotonicity constraint is replaced by $\mathbf{J}_{\phi}(y_{\mathrm{m},i})\succ 0$, where $\mathbf{J}_{\phi}(\cdot)$ denotes the Jacobian of $\phi(\cdot)$. The probit likelihood (28) is thus replaced by
$\mathrm{Pr}\left(\mathcal{M}_{i}|\mathbf{J}_{\phi}(y_{\mathrm{m},i})\right)\approx F_{\mathcal{N}}\left(\mathrm{det}\left(\mathbf{J}_{\phi}(y_{\mathrm{m},i})\right)/\nu_{0}\right).$ (31)
The virtual derivative points become the vectorized Jacobian:
$\Phi^{\prime}(\mathbf{y}_{\mathrm{m}})=\mathrm{col}\left(\mathrm{vec}\left(\mathbf{J}_{\phi}(y_{\mathrm{m},1})\right),\dots,\mathrm{vec}\left(\mathbf{J}_{\phi}(y_{\mathrm{m},n_{m}})\right)\right),$
and the definitions of $\mathbf{m}^{\prime}_{y}(\mathbf{y}_{\mathrm{m}})$, $\mathbf{k}^{\prime\prime}_{y}(\mathbf{y}_{\mathrm{m}},\mathbf{y}_{\mathrm{m}})$, $\kappa^{\prime}_{y}(\boldsymbol{\eta}^{d},\mathbf{y}_{\mathrm{m}})$ are modified accordingly. The EP algorithm remains unchanged except that 1) the approximate Gaussian likelihood becomes multivariate of dimension $n_{y}^{2}$, and 2) the closed-form solutions to the desired posterior marginal moments $\hat{\mu}_{i}$, $\hat{\sigma}_{i}^{2}$ (lines 6-8 in Algorithm 1) are no longer valid, cf. [35, Section 3.9], and numerical integration is required to calculate the moments.
From (62) and (63), the approximate posterior distribution with the virtual derivative points is given by:
$f(\eta)|\mathcal{M},\boldsymbol{\chi}^{d}\sim\mathcal{N}\left(m_{\mathrm{pm}}(\eta),k_{\mathrm{pm}}(\eta)\right),$ (32)
where
$m_{\mathrm{pm}}(\eta)$ $:=m(\eta)+\mathbf{k}^{\top}_{\mathrm{j}}(\eta)(K_{\mathrm{j}}+\tilde{\Sigma}_{\mathrm{j}})^{-1}\left(\tilde{\boldsymbol{\mu}}_{\mathrm{j}}-\mathbf{m}_{\mathrm{j}}\right),$ (33)
$k_{\mathrm{pm}}(\eta)$ $:=k(\eta,\eta)-\mathbf{k}^{\top}_{\mathrm{j}}(\eta)(K_{\mathrm{j}}+\tilde{\Sigma}_{\mathrm{j}})^{-1}\mathbf{k}_{\mathrm{j}}(\eta),$ (34)
Similar to (15), the hyperparameters can be estimated by maximizing the approximate marginal likelihood:
$q\left(\mathcal{M},\boldsymbol{\chi}^{d}|\Theta\right):=Z=(2\pi)^{-\frac{M+n_{m}}{2}}\mathrm{det}\left(\Lambda_{\mathrm{j}}(\Theta)\right)^{-\frac{1}{2}}\\ \exp\left(-\frac{1}{2}\delta_{\mathrm{j}}^{\top}\Lambda_{\mathrm{j}}^{-1}(\Theta)\delta_{\mathrm{j}}\right)\prod_{i=1}^{n_{m}}\tilde{Z}_{i},$ (35)
where we use $q(\cdot)$ to denote approximate probability densities after expectation propagation, $\delta_{\mathrm{j}}:=\tilde{\boldsymbol{\mu}}_{\mathrm{j}}-\mathbf{m}_{\mathrm{j}}$, and $\Lambda_{\mathrm{j}}(\Theta):=K_{\mathrm{j}}+\tilde{\Sigma}_{\mathrm{j}}$. Luckily, it was shown in [41, Appendix C.1.2] that the gradient of $\tilde{Z}_{i}$ with respect to the hyperparameters is zero. Thus, the hyperparameter estimation problem can be simplified as
$\Theta_{\mathrm{m}}$ $=\mathrm{arg}\underset{\Theta}{\mathrm{max}}\ q\left(\mathcal{M},\boldsymbol{\chi}^{d}|\Theta\right)$ (36)
$=\mathrm{arg}\underset{\Theta}{\mathrm{min}}\ \mathrm{logdet}\left(\Lambda_{\mathrm{j}}(\Theta)\right)+\delta_{\mathrm{j}}^{\top}\Lambda_{\mathrm{j}}^{-1}(\Theta)\delta_{\mathrm{j}}.$ (37)
Remark 7.
The approach presented in this section generalizes the results in [36] by allowing for non-zero prior mean functions and virtual derivative points on components of the nonlinear function rather than the whole function.
3.4 Posterior Prediction
@@
$\left[\Gamma_{1}\ \Gamma_{2}\right]\mathrm{col}\left(\Psi(\mathbf{u}),\Phi(\mathbf{y}_{\mathrm{p}})\right)-\Phi(\mathbf{y}_{\mathrm{f}})\sim\mathcal{N}\left(m_{\mathrm{pm}}(\eta),k_{\mathrm{pm}}(\eta)\right).$
@@
Let the true noise-free future output trajectory be $\mathbf{y}_{\mathrm{f,0}}$. From (2), we have $\left[\Gamma_{1}\ \Gamma_{2}\right]\mathrm{col}\left(\Psi(\mathbf{u}),\Phi(\mathbf{y}_{\mathrm{p}})\right)-\Phi(\mathbf{y}_{\mathrm{f,0}})=\left[\Gamma_{2}\ \mathbf{0}\right]\mathbf{e}$. Therefore, the GP model characterizes probable trajectories of (1) in the sense that for any given input trajectory $\mathbf{u}$ and past output trajectory $\mathbf{y}_{\mathrm{p}}$, $\mathbf{y}_{\mathrm{f}}$ is the future output trajectory with projected prediction error
$\Phi(\mathbf{y}_{\mathrm{f,0}})-\Phi(\mathbf{y}_{\mathrm{f}})\sim\mathcal{N}\left(m_{\mathrm{pm}}(\eta),\hat{k}_{\mathrm{pm}}(\eta)\right),$ (38)
where
$\hat{k}_{\mathrm{pm}}(\eta)=k_{\mathrm{pm}}(\eta)+\mathrm{cov}\left(\left[\Gamma_{2}\ \mathbf{0}\right]\mathbf{e}\right)=k_{\mathrm{pm}}(\eta)+\sigma^{2}\Gamma_{2}\Gamma_{2}^{\top}.$ (39)
The optimal posterior prediction of $\mathbf{y}_{\mathrm{f}}$ can be obtained by the maximum likelihood method. In detail, $\mathbf{y}_{\mathrm{f}}$ is estimated by maximizing the probability of observing $f(\eta)=\mathbf{0}$, i.e.,
$\mathbf{y}_{\mathrm{f}}$ $=\mathrm{arg}\underset{\mathbf{y}_{\mathrm{f}}}{\mathrm{max}}\ p\left(f(\eta)=\mathbf{0}|\mathcal{M},\boldsymbol{\chi}^{d}\right)$ (40)
$=\mathrm{arg}\underset{\mathbf{y}_{\mathrm{f}}}{\mathrm{min}}\ \mathrm{logdet}\left(k_{\mathrm{pm}}(\eta)\right)+m_{\mathrm{pm}}^{\top}(\eta)k_{\mathrm{pm}}^{-1}(\eta)m_{\mathrm{pm}}(\eta).$ (41)
Remark 8.
Other optimality criteria can also be used to obtain the optimal prediction depending on user preferences. This includes the minimum mean-squared error predictor:
$\mathbf{y}_{\mathrm{f}}=\mathrm{arg}\underset{\mathbf{y}_{\mathrm{f}}}{\mathrm{min}}\ \mathrm{tr}\left(\hat{k}_{\mathrm{pm}}(\eta)\right)+m_{\mathrm{pm}}^{\top}(\eta)m_{\mathrm{pm}}(\eta),$ (42)
and the minimum-variance unbiased predictor:
$\mathbf{y}_{\mathrm{f}}=\mathrm{arg}\underset{\mathbf{y}_{\mathrm{f}}}{\mathrm{min}}\ g\left(\hat{k}_{\mathrm{pm}}(\eta)\right)\ \mathrm{s.t.}\ m_{\mathrm{pm}}(\eta)=\mathbf{0},$ (43)
@@
where $g(\cdot)$ can be any optimality measure, such as trace, determinant, or the largest eigenvalue, i.e., the A-, D-, and E-optimality.
Remark 9.
@@
For Hammerstein systems, since $\mathbf{k}_{y}(\cdot,\cdot)=\mathbf{0}$ as mentioned in Remark 4, $k_{\mathrm{p}}(\eta)$ does not depend on $\mathbf{y}_{\mathrm{f}}$ and the monotonicity condition $\mathcal{M}$ is automatically satisfied. All optimality criteria admit their minimum when $m_{\mathrm{p}}(\eta)=\mathbf{0}$, i.e., $m(\eta)-\mathbf{k}^{\top}(\boldsymbol{\eta}^{d},\eta)\left(K+\bar{\Sigma}\right)^{-1}\mathbf{m}(\boldsymbol{\eta}^{d})=\mathbf{0}$. This leads to an explicit predictor by using (20):
$\mathbf{y}_{\mathrm{f}}=\left[\Gamma_{1}\ \Gamma_{2}\right]\mathrm{col}\left(\mathbf{m}_{u}(\mathbf{u}),\mathbf{y}_{\mathrm{p}}\right)-\mathbf{k}^{\top}(\boldsymbol{\eta}^{d},\eta)\left(K+\bar{\Sigma}\right)^{-1}\mathbf{m}(\boldsymbol{\eta}^{d}).$
3.5 Hyperparameter Tuning